OpenStudy (anonymous):

Find the area of a circle whose circumference is 12pi ft. A. 113.04 square feet B. 24 square feet C. 144 square feet D. 108.12 square feet

OpenStudy (anonymous):

help would be greatly appreciated

jimthompson5910 (jim_thompson5910):

you start by using the circumference formula C = 2*pi*r 12*pi = 2*pi*r ... plug in the given circumference 6 = r ... divide both sides by 2pi (to isolate r) r = 6 now you know the radius is 6 feet

jimthompson5910 (jim_thompson5910):

once you find the radius, you can find the area using A = pi*r^2

OpenStudy (anonymous):

im confused

jimthompson5910 (jim_thompson5910):

do you see how I got r = 6 ?

OpenStudy (anonymous):

yes but how do you multiply r^2

jimthompson5910 (jim_thompson5910):

r^2 means r squared

jimthompson5910 (jim_thompson5910):

so if r = 6, then r^2 = 6^2 = 6*6 = 36

jimthompson5910 (jim_thompson5910):

r = 6 r^2 = 36

OpenStudy (anonymous):

ok I know this ^ means square

jimthompson5910 (jim_thompson5910):

^ means exponent actually

jimthompson5910 (jim_thompson5910):

when you have a 2 after the ^, it means square ^2 means squared

jimthompson5910 (jim_thompson5910):

^3 means cubed etc etc

OpenStudy (anonymous):

so the answers a

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks for your help

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Find the area of a circle whose circumference is 12pi ft.